9. A car moving initially with a speed of S m/s brakes to a stop in 10 s. The average braking force is 700 N. Find mass of the car.
Answer #1

(9) Initial speed of the car, u = 5.0 m/s

The car stop in time, t = 10.0 s

Final speed of the car, v = 0

Suppose, a = acceleration of the car

Use the expression –

v= u + a*t

put the values –

0 = 5.0 m/s + a*10.0 s

=> a = -5/10 = -0.50 m/s^2

So, magnitude of the acceleration, a = 0.50 m/s^2

Negative sign shows that the car is slowing down.

Again. suppose, m = mass of the car

Average force, F = 700 N (given)

Now –

F = m*a

=> m = F/a = 700 N / (0.50 m/s^2) = 1400.0 kg

Hence, mass of the car = 1400.0 kg (Answer)